OpenStudy (anonymous):

Solve for x. fourteen x plus twelve equals one hundred ten x equals negative seven x equals seven x equals eight x equals nine

OpenStudy (solomonzelman):

fourteen x plus twelve equals one hundred ten ---> 14x+12=110 subtract 12 from both sides 14x+12-12=110-12 14x=98 divide both sides by 14 14x/14=98/14 x=7

OpenStudy (solomonzelman):

So it is B, x equals seven